OpenStudy (anonymous):

population of rabbits is given by p(t)=500/1+e^-t determine the rate of growth after 3 years

OpenStudy (amistre64):

rate of growth sounds like its asking for a derivative at 3 years

OpenStudy (anonymous):

why is that e in the function. i dont know how to get the answer because of the e

OpenStudy (amistre64):

the "e" represents a transmutational number ... transmigratory? trans-something ... its a number that comes up alot in natural phenomenae

OpenStudy (amistre64):

p(t)=500/(1+e^-t) p(t) = 500 (1+e^-t)^-1 p'(t) = 500 -(1+e^-t)^-2 (-e^-t) = (-500 e^-t)/(1+e^-t)^2 p'(3) = (-500 e^-3)/(1+e^-3)^2 =abt -22.588.... this tells us the growth rate is declining by about 22.588 after 4 years

OpenStudy (anonymous):

but what is the valu of e?

OpenStudy (amistre64):

it has a "value" attributed to it just like "pi" does. e = 2.71828182845905....... if i recall it correctly